Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

When tossing 5 coins simultaneously, find the probability that at least 1 head is showing.

OpenStudy (anonymous):

The probability that a particular coin is showing tails is .5. The probability that two coins are BOTH showing tails is .5 X .5 = .25 (reducing probability with each coin added). The probability of no coin showing heads is .5 X .5 X .5 X .5 X .5 = (.5)^5 = .03125. Therefore, the probability of at least 1 coin showing heads is 1 - .03125 = .96875

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!